OpenStudy (anonymous):

Create your own radical expression, containing a numerical coefficient and two different variables, which can be simplified. Explain, in complete sentences, how you developed the expression, the process of simplifying it, and provide the final simplified expression.

OpenStudy (anonymous):

Let\[a,b,c\in\mathbb{R}.\]Then\[a\sqrt{b^2c}=a\sqrt{b^2}\sqrt{c}=ab\sqrt{c}.\]

OpenStudy (anonymous):

You can also say\[\sqrt{d}=\sqrt{cb}=\sqrt{a^2b}=a\sqrt{b}.\]